# A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective

### v1

@@ D.1 Generated Images over Iterations
D.2 Generated Image Distribution Becomes Spiky
@@
We want to clarify the nuance between MNND and variance. A spiky distribution does not imply that all data points are concentrated in a single small region—this behavior has already been illustrated by the variance collapse behavior [Shumailov et al., 2024, Kazdan et al., 2024, Bertrand et al., 2024]. Specifically, Figure˜18 shows that the variance of the generated dataset only slightly decreases along the successive iterations and is far from complete collapse. On the contrary, MNND decreases almost exponentially and reaches a small value after $10$ iterations. Thus, the results align with the prior claim in Schaeffer et al. [2025], Kazdan et al. [2024] that the collapse of variance progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the collapse in MNND emerges at an early stage, highlighting a critical memorization issue caused by training on synthetic data.
@@
D.3 Self-consuming Loop with Fresh New Data
@@
D.4 Additional Metric
Since our Greedy Section method is performed on the feature space extracted by the DINO model, we also use FDDINO [Stein et al., 2023] as an alternate for FID to measure the quality of the generated images. As we show in Figure˜19, the Greedy Selection method can also improve the FDDINO metric compared to the vanilla paradigms.
@@ D.4 Additional Metric
D.5 Additional Results for Section˜5
We provide visualization of the generated FFHQ images with their nearest training neighbors to show the model’s generalizability. Figure˜20 shows some images for different training paradigms at $5$-th iteration in grid format. With augmentation from the Greedy Selection method, the model generates images that deviate more from the training set compared to the vanilla accumulate paradigm, thereby enhancing its generalization ability.

### v2

@@ D.3 Generated Images over Iterations
D.4 Generated Image Distribution Becomes Spiky
@@
We want to clarify the nuance between MNND and variance. A spiky distribution does not imply that all data points are concentrated in a single small region—this behavior has already been illustrated by the variance collapse behavior [Shumailov et al., 2024, Kazdan et al., 2024, Bertrand et al., 2024]. Specifically, Figure˜20 shows that the variance of the generated dataset only slightly decreases along the successive iterations and is far from complete collapse. On the contrary, MNND decreases almost exponentially and reaches a small value after $10$ iterations. Thus, the results align with the prior claim in Schaeffer et al. [2025], Kazdan et al. [2024] that the collapse of variance progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the collapse in MNND emerges at an early stage, highlighting a critical memorization issue caused by training on synthetic data.
@@
D.5 Self-consuming Loop with Fresh New Data
@@
D.6 Additional Metric
Since our Greedy Section method is performed on the feature space extracted by the DINO model, we also use FDDINO [Stein et al., 2023] as an alternate for FID to measure the quality of the generated images. As we show in Figure˜21, the Greedy Selection method can also improve the FDDINO metric compared to the vanilla paradigms.
@@ D.6 Additional Metric
D.7 Visualization of the Improvement on Generalization Score
We provide visualization of the generated FFHQ images with their nearest training neighbors to show the model’s generalizability. Figure˜22 shows some images for different training paradigms at $5$-th iteration in grid format. With augmentation from the Greedy Selection method, the model generates images that deviate more from the training set compared to the vanilla accumulate paradigm, thereby enhancing its generalization ability.